Rock It Out had the following account balances on January 1:

Cash: $836,000

Raw Materials: $18,500

WIP: $115,400

Finished Goods Inventory $45,600

The following took place during the year:

1. Purchased raw material with cash $173,600.

2. Put material into production:

$129,000 of direct material

$12,000 of indirect material.

3. Paid payroll:

$158,000 in wages to direct labor

$93,000 salary for supervisors

4. Applied OH based on a pre-determined rate of $13 per machine hour. Actual machine hours were 9,950.

5. Incurred and paid other overhead items of $35,000.

6. Transferred items costing $473,500 to finished goods.

7. Sold goods costing $453,900 for cash of $886,200.

Assume that the Raw Material Inventory account contains both direct and indirect material. The company uses normal costing.

  1. What is the ending balance of the raw materials account for Dec 31?
  2. What is the ending balance of the WIP account for Dec 31?
  3. What is the ending balance of the Finished Goods account for Dec 31?
  4. How much is revenue?
  5. How much is COGS before closing the MOH account?
  6. How much was over- or under-applied OH?
Answer #1

a) Ending balance of raw material = 18500+ 173600 - 129000-12000 = 51100

b) Ending balance of WIP account = (115400 + 129000 + 158000+ 129350 ) -473500 = $58250

c) Ending balance of Finished goods account = 45600+ 473500-453900 = $65200

d) Revenue as mentioned is 886200

e)Actual overhead = 12000+93000+35000 = 140000

Applied overhead = 129350

Under applied overhead = 140000-129350 = $10650